# IEEEexample:BSTcontrol

### v1

@@ 3 Experiments
Silhouette
1.364
1.136 1.091 1.197
Eigengap
1.318
1.455
1.636
1.470
DB 1.091
1.818
1.500
1.470
CH
1.364
1.318 1.227
1.303
Random
3.909
2.455
3.091
3.152
Voting 1.091
1.455
1.046 1.197
Average 1.091
1.364
1.091 1.182
@@
Eigengap
45.45
59.09 40.91
48.48
DB
54.55
31.82
40.91
42.42
CH 54.55
31.82
68.18
51.52
Random
31.82
31.82
27.27
30.30
Voting
54.55
40.91
59.09
51.52
Average 54.55
45.45
63.64 54.54
@@
Silhouette
15.99
19.68 12.78 16.16
Eigengap
16.36
25.01
16.47 19.28
DB 14.70
26.16
14.11 18.32
CH
18.03
26.88
12.09 19.01
Random
27.05
26.08
21.54 24.89
Voting
15.06
24.01
12.04 17.04
@@
Silhouette
0.00
1.75
3.20
1.00
Eigengap
0.33
0.75
4.0
9.00
DB
1.167
3.25
1.00
1.00
CH
0.75
1.50
2.40
0.00
Voting
0.00
1.50
3.20
1.00
Average
0.00
1.75
3.20
1.00
Avg. MSE
0.375
1.75
2.83
2.17
@@
Silhouette
6.10
20.03
23.43
10.52
Eigengap
10.74
24.09
22.51
24.61
DB
10.40
20.44
18.67
10.52
CH
7.95
17.72
17.75
10.96
Voting
6.10
18.21
21.67
10.52
Average
6.10
18.16
21.67
10.52
Avg. Error
7.90
19.78
20.95
12.94
Baseline
3.31
8.23
13.75
6.04

### v2

@@ 3 Experiments
Silhouette 1.364 1.136 1.091 1.197
Eigengap 1.318 1.455 1.636 1.470
DB 1.091 1.818 1.500 1.470
CH 1.364 1.318 1.227 1.303
Random 3.909 2.455 3.091 3.152
Voting 1.091 1.455 1.046 1.197
Average 1.091 1.364 1.091 1.182
@@
Eigengap 45.45 59.09 40.91 48.48
DB 54.55 31.82 40.91 42.42
CH 54.55 31.82 68.18 51.52
Random 31.82 31.82 27.27 30.30
Voting 54.55 40.91 59.09 51.52
Average 54.55 45.45 63.64 54.54
@@
Silhouette 15.99 19.68 12.78 16.16
Eigengap 16.36 25.01 16.47 19.28
DB 14.70 26.16 14.11 18.32
CH 18.03 26.88 12.09 19.01
Random 27.05 26.08 21.54 24.89
Voting 15.06 24.01 12.04 17.04
@@
Silhouette 0.00 1.75 3.20 1.00
Eigengap 0.33 0.75 4.0 9.00
DB 1.167 3.25 1.00 1.00
CH 0.75 1.50 2.40 0.00
Voting 0.00 1.50 3.20 1.00
Average 0.00 1.75 3.20 1.00
Avg. MSE 0.375 1.75 2.83 2.17
@@
Silhouette 6.10 20.03 23.43 10.52
Eigengap 10.74 24.09 22.51 24.61
DB 10.40 20.44 18.67 10.52
CH 7.95 17.72 17.75 10.96
Voting 6.10 18.21 21.67 10.52
Average 6.10 18.16 21.67 10.52
Avg. Error 7.90 19.78 20.95 12.94
Baseline 3.31 8.23 13.75 6.04